# Signs as Tokens: An Autoregressive Multilingual Sign Language Generator

### v1

@@ 3.4 Autoregressive Multilingual Generator
Figure 4: Part-wise decoding. To enable motion token decoding for different body parts, we develop a part-wise decoding strategy. This approach ensures that the decoding process always begins with a special token, such as <$\mathtt{Lang\_B}$>for the upper-body motion sequence and <$\mathtt{Lang\_LH}$>for the left-hand motion sequence. The placeholder $\mathtt{Lang}$ should be replaced with the target sign language, such as $\mathtt{CSL}$ and $\mathtt{ASL}$ for Chinese and American sign languages, respectively.
Using obtained discrete sign representations, we can approach SLG similarly to natural language processing with a pretrained Transformer-based LM [32]. The proposed approach enables more robust SLG across diverse discourse domains and supports multiple sign languages within a single unified model.
Integrated Vocabulary. We first gather the discrete tokens for different body parts to construct a part-wise motion vocabulary: $\mathcal{V}_{m}^{p}=\{\textless\mathtt{p\_i}\textgreater\}_{i=1}^{N_{Z}^{p}}$, e.g., <$\textless\mathtt{B\_1}\textgreater$ > and <$\textless\mathtt{RH\_2}\textgreater$ >. The overall motion vocabulary is a combination of three part-wise vocabularies: $\mathcal{V}_{m}=\{\mathcal{V}_{m}^{B},\mathcal{V}_{m}^{LH},\mathcal{V}_{m}^{RH}\}$.
@@
Part-wise Decoding. Since we decompose the motion sequence into three parts via the decoupled tokenizer, we accordingly devise a part-wise decoding strategy such that the LM can output token sequences for each body part. As shown in Figure 4, given a text prompt $\mathbf{X}=\{x_{i}\}_{i=1}^{L}$, where $x_{i}\in\mathcal{V}_{t}$, the LM encoder first extracts latent features, $\mathbf{E}\in\mathbb{R}^{T\times C_{e}}$, where $C_{e}$ denotes the number of dimensions, serving as the condition for the following decoding process.
Moreover, we instantiate three LM decoders with shared weights, each one in charge of a single body part. To prompt the decoders with the information of body parts, decoding process starts from one of the above mentioned special tokens, based on the target language and body part. For example, the starting token would be set to <$\mathtt{ASL\_B}$>when the model is required to generate upper-body motions for ASL. Then in each decoding step, the decoders will predict the probability distributions of the next token $m_{k}^{p}\in\mathcal{V}_{m}^{p}$, conditioned on the encoder hidden states $\mathbf{E}$ and the previous predictions: $q(m_{k}^{p}|\mathbf{E})=\prod_{k}q(m_{k}^{p}|m_{<k}^{p},\mathbf{E})$. To train the LM, we utilize the traditional cross-entropy loss:
$\mathcal{L}_{LM}^{p}=-\sum_{k=1}^{K}\log q(m_{k}^{p}|m_{<k}^{p},\mathbf{E}),$ (2)
where $K$ denotes the length of the target motion sequence.
@@
Similar to the tokenizer training scheme, the overall loss function for the LM training stage, $\mathcal{L}_{LM}$, is also a sum over all body parts. During inference, we adopt a simple greedy decoding algorithm that the model always outputs the token with the highest probability at each step. The obtained token sequences will be fed into the corresponding part-decoder, $\mathcal{D}_{p}(\cdot)$, to reconstruct sign motions.
@@ 4 Experiments
Datasets. We construct a multilingual sign language dataset by merging two widely adopted datasets: How2Sign [8] and CSL-Daily [79]. How2Sign is a large-scale American sign language dataset, consisting of 31K, 1.7K, and 2.3K video-text pairs in training, dev, and test sets, respectively. CSL-Daily is the largest existing continuous Chinese sign language dataset, containing 18K samples for training, 1K for development, and 1.1K for testing. To train SOKE, we use the SMPL-X annotations of How2Sign dataset as provided in [3] along with the curated SMPL-X annotations of CSL-Daily as discussed in Section 3.1.
Evaluation Metrics. To quantitatively assess the reconstruction quality of our proposed DETO, we use Procrustes-aligned mean per vertex position error (PA-MPVPE) and mean per joint position error (PA-MPJPE) as described in [31, 69]. Additionally, since the length of generated signs may differ from the ground truth, we employ the well-established dynamic time warping (DTW) [37] on joint errors, which calculates sequence-level distances between the generated signs and ground truth. Furthermore, in line with [3], we also report back-translation scores, including BLEU-4 [39] and ROUGE [30], to assess the interpretability of the generated signs.
Figure 5: Qualitative comparisons of reconstructed signs on CSL-Daily (first two columns) and How2Sign (last two columns). The results verify that using our proposed decoupled tokenizer (DETO) significantly enhances the reconstruction quality of hands.
(a) CSL-Daily.
(b) How2Sign.
Figure 6: Qualitative comparisons of generated signs between our proposed method, SOKE, with the baseline method, MotionGPT [22], on CSL-Daily and How2Sign. The results indicate that SOKE can generate finer-grained hand articulations than MotionGPT.
Implementation Details. For DETO, we empirically set the code numbers to $N_{Z}^{B}=96$ for the body and $N_{Z}^{LH}=N_{Z}^{RH}=192$ for the left and right hands, respectively, with a code dimension of $C=512$. Following [22], the motion encoders consist of three 1D-CNN blocks with a temporal downsampling factor of 4. We train DETO with a batch size of 256 per GPU for 500 epochs, using the AdamW optimizer [33] and a cosine learning rate scheduler starting at 2e-4. In line with [22], we employ EMA codebook reset techniques [46] to enhance the utilization rate of codebooks. For the multilingual language model, we use mBART-large-cc25111https://huggingface.co/facebook/mbart-large-cc25 [32], which features 12 encoder-decoder layers with a model dimension of 1024 and is widely used in various sign language translation methods [6, 62, 5, 84, 24]. The language model is fine-tuned with a batch size of 24 per GPU for 150 epochs, employing the same optimizer settings as DETO. All models are trained using 6 RTX 3090 GPUs.
Method Multi- lingual How2Sign CSL-Daily
@@
Body L-Hand R-Hand BLEU-4 ROUGE Body L-Hand R-Hand BLEU-4 ROUGE
Prog. Trans. [50] × 14.15 11.21 11.93 2.75 29.87 15.98 12.54 13.28 3.07 28.44
Adv. Train. [51] × 13.78 11.02 11.32 6.21 32.33
NAR [21] × 13.94 11.33 12.26 5.75 31.98
Text2Mesh [58] × 13.99 11.59 12.29 7.51 33.17 13.47 12.04 12.16 5.11 31.62
MotionGPT [22] × 11.23 4.23 4.55 11.45 43.87 10.81 3.84 3.71 8.82 38.26
NSA [3] × 7.83 6.74 7.91 13.12 47.55
SOKE (ours) 7.61 2.83 2.93 14.58 49.64 8.08 2.20 2.66 11.90 43.08
Table 1: Comparison with SOTA SLG methods. denotes reimplementation results of open-source methods.
Method Multi- lingual How2Sign CSL-Daily
@@
All Hand Face Body Hand All Hand Face Body Hand
SignVAE [69] × 32.23 10.16 1.70 26.72 10.53 39.44 8.83 1.59 32.89 9.15
MotionVAE [22] × 29.49 8.18 1.69 25.47 8.42 32.17 6.29 1.39 27.13 6.48
DETO (ours) × 24.08 7.08 1.26 19.73 7.28 30.72 5.65 1.21 25.36 5.78
SignVAE [69] 30.19 9.85 1.55 24.90 10.21 38.06 8.42 1.54 31.81 8.71
MotionVAE [22] 25.73 7.91 1.45 22.08 8.18 31.75 6.20 1.44 27.03 6.38
DETO (ours) 23.26 6.47 1.11 19.00 6.65 28.10 5.10 1.17 23.27 5.20
Table 2: Comparison with SOTA sign motion tokenizers. denotes reimplementation results.
4.1 Qualitative Evaluation
Reconstructed Signs. As depicted in Figure 5, we investigate the impact of decoupled tokenizer (DETO) on sign reconstruction. The visualizations reveal that substituting DETO with a single VQ-VAE for full-body motion reconstruction leads to diminished quality in hand movements, especially when the signer is performing complex handshapes (the first and second columns of Figure 5).
Generated Signs. We first reimplement a state-of-the-art (SOTA) motion generation approach, MotionGPT [22], which utilizes a single VQ-VAE (MotionVAE) to discretize whole-body motions and employs a monolingual language model [45], on sign language datasets. In Figure 6, we perform a qualitative comparison between our proposed method, SOKE, and MotionGPT. As can be easily seen, our proposed method produces significantly better results, with visual quality comparable to the ground truth sign videos in both the ASL and CSL datasets. Notably, our method can generate precise and informative hand movements, whereas MotionGPT struggles with poor hand motion generation, suggesting the effectiveness of decoupled tokenizing for hands. Additionally, the proposed method supports both sign languages simultaneously, whereas MotionGPT is monolingual and requires separate models for each sign language dataset.
4.2 Quantitative Evaluation
Comparison with SOTA SLG Methods. In addition to the above qualitative comparison, we also perform a quantitative evaluation against SOTA SLG methods, as detailed in Table 1. Text2Mesh [58] is a pioneering work in generating 3D sign motions from texts. However, the model depends on glosses as an intermediate supervision to learn sign-text mappings. This reliance limits its ability to generalize to large-scale datasets that lack gloss annotations. Additionally, glosses have inherent limitations, such as inadequate representation [38], which can cause significant error propagation, resulting in high DTW errors (>11 for both hands). Neural Sign Actors (NSA) [3] currently represents the leading SLG approach, framing SLG as a motion generation task and utilizing a diffusion model as its backbone. However, NSA is a monolingual generator, tested solely on How2Sign, which raises concerns about its ability to generalize across different sign languages. In comparison, our method achieves a similar DTW error for body and significantly lower errors for both hands, demonstrating the effectiveness of our decoupled tokenizer. Furthermore, with just a single unified model, our approach achieves very low hand errors of 2.20/2.66 on CSL-Daily.
Setting How2Sign CSL-Daily
@@ 4.2 Quantitative Evaluation
All Hand Face Body Hand All Hand Face Body Hand
Whole-Body 25.73 7.91 1.45 22.08 8.18 31.75 6.20 1.44 27.03 6.38
Body (w/ face) + Hands 23.63 7.59 1.13 19.34 7.82 28.34 5.85 1.18 23.34 5.99
Body (w/ face) + L-Hand + R-Hand (default) 23.26 6.47 1.11 19.00 6.65 28.10 5.10 1.17 23.27 5.20
Body + L-Hand + R-Hand + Face 24.01 6.49 1.07 19.97 6.68 28.55 5.18 1.18 23.56 5.23
Table 3: Ablation study on the proposed decoupled tokenizer.
Comparison with SOTA Sign Motion Tokenizers. To demonstrate the effectiveness of our decoupled tokenizer (DETO), we reimplement two SOTA sign motion tokenizers: SignVAE [69] and MotionVAE [22]. Both approaches employ a single VQ-VAE to tokenize full-body motions, while MotionVAE incorporates advanced training techniques such as EMA reset [46] to enhance codebook utilization. As illustrated in Table 2, our DETO significantly surpasses both methods across all metrics, including reconstruction errors for various body parts, by a substantial margin. These results strongly indicate the benefits of decoupled modeling for sign languages. Additionally, Table 2 reveals that tokenizer training benefits from multilingual data; all methods show consistent performance improvements when trained on combined datasets compared to training on each dataset individually. This is consistent with the finding in [62], in which the authors leverage cross-lingual signs to boost sign language recognition.
Ablation Study on Decoupled Tokenizer. Inspired by the multi-cue characteristics of sign languages, we introduce a decoupled tokenizer that discretizes sign motion sequences for different body parts. In Table 3, we present an ablation study on various configurations: a single VQ-VAE for whole-body motions, one VAE for the body and another for both hands (Body+Hands), and an additional VAE for facial expressions. The findings indicate that progressively separating the body and hands leads to significantly improved results, particularly benefiting hand reconstruction, with hand joint errors decreasing from 7.82 to 6.65 on How2Sign and from 5.99 to 5.20 on CSL-Daily. However, we did not observe clear performance improvements when adding a VAE for facial expressions, likely due to the low dimensionality of facial expression parameters, which are limited to 10 dimensions. In the future, we aim to employ more advanced facial expression models to annotate detailed expression parameters for sign language datasets.
@@
Body L-Hand R-Hand Body L-Hand R-Hand
12.91 3.41 3.71 11.58 3.56 3.13
11.08 3.08 3.38
9.27 3.23 3.22
7.61 2.83 2.93 8.08 2.20 2.66
Table 4: Study on the impact of decoupled tokenizer (DETO), How2Sign (H2S) data, and CSL-Daily data on SLG performance.
Impact of Tokenizer and Multilingual Data on SLG. In Table 4, we examine the effects of using a decoupled tokenizer and multilingual data on the overall generation performance. The results suggest two key insights. First, DETO is crucial for achieving high-quality generation performance, as its removal results in an increase of 1.36 in hand DTW errors on CSL-Daily, underscoring the effectiveness of decoupled modeling in sign language generation. Second, incorporating multilingual data enhances generation performance, primarily due to the development of a more robust tokenizer, as highlighted in Table 2.

### v3

@@ 3.3 Autoregressive Multilingual Generator
Using obtained discrete sign representations, we can respect the discrete properties of sign languages and approach SLG similarly to text generation with a pretrained LM [40].
@@
Sequential Decoding. Following previous motion generation works [42, 74], a naive decoding method is simply flattening tokens from each body part into a single sequence with tripled length: $\mathbf{Y}=\{y_{1}^{B},y_{1}^{LH},y_{1}^{RH},\dots,y_{K}^{B},y_{K}^{LH},y_{K}^{RH}\}$, where $y_{k}^{p}\in\mathcal{V}_{m}^{p}$ denotes the part-wise motion token, and $K$ denotes the length of the target sign motion sequence. The decoding process can then be formulated as:
$P(\mathbf{Y}|\mathbf{h}_{en})=\prod_{k=1}^{K}P(y_{k}^{B,LH,RH}|y_{<k}^{B,LH,RH})$ (2)
$=\prod_{k=1}^{K}P(y_{k}^{B}|y_{<k}^{\{p\}})P(y_{k}^{LH}|y_{k}^{B},y_{<k}^{\{p\}})P(y_{k}^{RH}|y_{k}^{B,LH},y_{<k}^{\{p\}}),$
where $\mathbf{h}_{en}$ denotes the last encoder hidden states. For simplicity, we write $y_{<k}^{B,LH,RH}$ as $y_{<k}^{\{p\}}$ and omit $\mathbf{h}_{en}$. We term this decoding method as sequential decoding (Figure 3(a)), which results in inefficient inference process requiring $3K$ decoding steps.
Parallel Decoding. In real-world applications, the model’s real-time capability is important. To accelerate the decoding process, a straightforward way is to decompose the target sequence $\mathbf{Y}$ into three part-wise sequences: $\mathbf{Y}^{p}=\{y_{1}^{p},\dots,y_{K}^{p}\}$. Moreover, three LM decoders are instantiated with shared weights, each one in charge of a single body part. To prompt the decoders with the information of body parts, we replace the language identifier tokens with a new set of special tokens: $\mathcal{V}_{l}=\{\textless\mathtt{Lang\_p}\textgreater\}$ > }, where $\mathtt{Lang}$ and $\mathtt{p}$ denote the target language and body part, respectively.
As shown in Figure 3(b), three parallel decoding processes start from one of the above mentioned special tokens, based on the target language and body part. For example, the starting token would be set to <$\mathtt{ASL\_B}$>when the model is required to generate upper-body motions for ASL. The decoding process can then be formulated as:
$P(\mathbf{Y}|\mathbf{h}_{en})=P(\mathbf{Y}^{B})P(\mathbf{Y}^{LH})P(\mathbf{Y}^{RH})$ (3)
$=\prod_{k=1}^{K}P(y_{k}^{B}|y_{<k}^{B})\prod_{k=1}^{K}P(y_{k}^{LH}|y_{<k}^{LH})\prod_{k=1}^{K}P(y_{k}^{RH}|y_{<k}^{RH}).$
Comparing Eq. 3 with Eq. 2, the parallel decoding strategy significantly reduces the dependence on previous tokens, as the decoding processes for each body part are completely separated. While this formulation enables efficient parallel execution, the overly independent assumption may result in suboptimal performance.
Multi-Head Decoding. To balance efficiency and information fusion across body parts, we propose a multi-head decoding strategy. As shown in Figure 3(c), we design three language modeling heads, implemented as fully connected layers, to predict motion tokens for each body part simultaneously at each step. Moreover, the decoder inputs at each step are modified to a weighted average of the token embeddings from different body parts. Specifically, denoting the token embeddings from upper body, left hand, and right hand as $\mathbf{E}^{B}$, $\mathbf{E}^{LH}$, and $\mathbf{E}^{RH}$, respectively, the input embeddings are defined as: $\mathbf{E}=(1-2\lambda)\mathbf{E}^{B}+\lambda\mathbf{E}^{LH}+\lambda\mathbf{E}^{RH}$, where $\lambda\in(0,0.5)$ is a hyper-parameter to control the weight of hand embeddings. The decoding process can then be formulated as:
$P(\mathbf{Y}|\mathbf{h}_{en})=\prod_{k=1}^{K}P(y_{k}^{B,LH,RH}|y_{<k}^{B,LH,RH})$ (4)
$=\prod_{k=1}^{K}P(y_{k}^{B}|y_{<k}^{\{p\}})P(y_{k}^{LH}|y_{<k}^{\{p\}})P(y_{k}^{RH}|y_{<k}^{\{p\}}),$
which can be viewed as relaxing Eq. 2 under a conditional independence assumption.
@@
Training and Inference. The LM is trained with a standard cross-entropy loss: $\mathcal{L}_{LM}=-\log P(\mathbf{Y}|\mathbf{h_{en}})$. During inference, we adopt a simple greedy decoding algorithm that the LM always outputs the token with the highest probability at each step. For parallel and multi-head decoding, the inference process terminates as soon as any head or decoder predicts an end-of-sentence token. The obtained token sequences will be fed into the corresponding part-decoder, $\mathcal{D}_{p}(\cdot)$, to reconstruct sign motions.
3.4 Retrieval-Enhanced SLG
Figure 4: Illustration of retrieval-enhanced SLG. We utilize external sign dictionaries to retrieve accurate word-level signs, which serve as additional conditions to enhance the generation quality.
Inspired by the effectiveness of retrieval-augmented generation in NLP [5, 12, 68], we propose to leverage external sign dictionaries to improve the fidelity of generated signs.
As shown in Figure 4, we first build a word-level sign dictionary for each target sign language by utilizing resources from isolated sign language recognition datasets [32, 37, 94, 85, 35] and online sources. Through our pose fitting pipeline, each RGB video in these dictionaries is converted into SMPL-X poses. Next, we employ a well-trained decoupled tokenizer to map the poses into discrete tokens. The resulting dictionary is represented as a set of quadruples $\{(w,m^{B},m^{LH},m^{RH})\}$, where $w$ denotes lemmatized words in the written language, and $m^{p}$ denotes motion tokens. If a word has multiple sign instances in the dictionary, we retain only the instance with the lowest reconstruction error after passing through the tokenizer.
Given a text input $\mathbf{X}=\{x_{i}\}_{i=1}^{L}$, consisting of $L$ words [^1] Since Chinese sentences lack spaces between words, we use Jieba [17, 45] for word segmentation., we collect all words and corresponding tokens found in the dictionary. The final prompt is created by concatenating raw texts $\mathbf{X}$ with the retrieved motion tokens $\{m_{j}^{p}\}$, which is then fed into the LM encoder to generate hidden states $\mathbf{h}_{en}$. Several examples are shown in Figure 4.
Figure 5: Qualitative comparisons of generated signs between our proposed method, SOKE, with the SOTA method, S-MotionGPT [27], on the test sets of How2Sign (left), CSL-Daily (middle), and Phoenix-2014T (right).
@@ 4 Experiments
Datasets. We construct a multilingual sign language dataset by merging three widely-adopted datasets: How2Sign [11], CSL-Daily [89], and Phoenix-2014T [4]. How2Sign is a large-scale ASL dataset, consisting of 35K video-text pairs. CSL-Daily is the largest existing continuous CSL dataset, containing 20K samples. Phoenix-2014T is a DGS dataset with 8K samples. To train SOKE, we use the SMPL-X poses of How2Sign dataset as provided in [3] along with the curated poses of CSL-Daily and Phoenix-2014T.
Evaluation Metrics. Since the length of generated signs may differ from the ground truth, we employ the well-established dynamic time warping (DTW) [3, 2] on both procrustes-aligned and original joint position errors (JPE) to measure sequence-level distances between the generated signs and ground truth. Following [3], we also report back-translation (B-T) scores in BLEU-4 [46] to assess the interpretability of the generated signs.
Implementation Details. For DETO, we empirically set the code numbers to $N_{Z}^{B}=96$ for the body and $N_{Z}^{LH}=N_{Z}^{RH}=192$ for the left and right hands, respectively, with a code dimension of $C=512$. We train DETO on the combined multilingual sign language dataset and sign dictionaries with a batch size of 320 per GPU for 500 epochs, using the AdamW optimizer [41] and a cosine learning rate scheduler starting at 2e-4. For the LM, we use mBART-large-cc25 [40], which features 12 encoder-decoder layers with a model dimension of 1024 and is widely used in various sign language translation methods [8, 69, 7, 96, 30]. We empirically set $\lambda=1/3$ to balance part-wise token embeddings. The LM is fine-tuned with a batch size of 32 per GPU on the multilingual sign language dataset for 150 epochs, employing the same optimizer settings as DETO. All models are trained using 6 RTX 3090 GPUs.
Method Multi- lingual How2Sign CSL-Daily Phoenix-2014T
@@
Body Hand Body Hand BLEU-4 Body Hand Body Hand BLEU-4 Body Hand Body Hand BLEU-4
NAR [26] × 13.94 11.80 5.75
Prog. Trans. [53] × 14.15 11.57 14.74 30.17 2.75 15.98 12.91 16.30 32.63 3.07 13.67 11.95 15.01 31.77 4.94
Text2Mesh [60] × 13.99 13.47 15.50 32.97 7.51 13.47 12.10 13.76 30.37 5.11 13.48 12.06 14.04 31.64 5.81
Adv. Train. [54] × 13.78 11.17 6.21
T2S-GPT [76] × 11.48 6.39 12.65 18.44 11.20 11.94 5.93 12.32 15.43 8.94 10.38 6.47 11.65 19.09 9.06
NSA [3] × 7.83 7.33 13.12
S-MotionGPT [27] × 11.23 4.39 12.41 13.74 11.45 10.81 3.78 11.58 11.31 8.82 9.45 3.41 10.42 9.08 9.68
SOKE (ours) 6.82 2.35 7.75 10.08 14.48 6.24 1.71 7.38 9.68 11.30 4.77 1.38 6.04 7.72 11.87
Table 1: Comparison with state-of-the-art sign language generation (text-to-sign) methods. denotes reimplementation results.
Decoding Method Sign Retrieval How2Sign CSL-Daily Phoenix-2014T
@@
Avg Body Hand s/video Avg Body Hand s/video Avg Body Hand s/video
Sequential × 4.64 8.55 3.53 3.26 3.74 8.14 2.49 3.15 3.26 6.97 2.20 3.23
Parallel × 5.06 9.27 3.86 1.39 4.19 8.61 2.93 1.41 3.54 7.02 2.55 1.35
Multi-head × 4.17 7.91 3.10 1.46 3.37 7.58 2.17 1.45 2.80 6.16 1.85 1.44
Multi-head 3.34 6.82 2.35 1.55 2.72 6.24 1.71 1.52 2.13 4.77 1.38 1.51
Table 2: Ablation study for the decoding method and sign retrieval. Latencies are measured using a single RTX 3090 GPU.
4.1 Comparison with State-of-the-Art Methods
Qualitative Comparison. We first reimplement a state-of-the-art (SOTA) open-sourced motion generation approach, MotionGPT [27], which utilizes a VQ-VAE to discretize whole-body motions and employs a monolingual language model [50], on sign language datasets. The resulting model, S-MotionGPT, achieves competitive results on sign language benchmarks (Table 1). In Figure 5, we perform a qualitative comparison between our SOKE and S-MotionGPT. As can be easily seen, our proposed method produces significantly better results, with visual quality comparable to the ground truth sign videos across the ASL, CSL, and DGS datasets. Notably, our method can generate precise and informative hand movements, which play a key role in conveying semantics in sign languages. Additionally, the proposed method supports three sign languages simultaneously, whereas S-MotionGPT is monolingual and requires separate models for each sign language dataset.
Quantitative Comparison. In addition to the above qualitative comparison, we also perform a quantitative evaluation against SOTA SLG methods, as detailed in Table 1. Our reimplemented S-MotionGPT achieves competitive results, with low DTW errors of 4.39 and 3.78 on How2Sign and CSL-Daily, respectively, demonstrating the effectiveness of formulating SLG as a language modeling task. The previous best SLG approach, Neural Sign Actors (NSA) [3], formulates SLG as a motion generation task and employs a diffusion model as its backbone. However, this approach overlooks the linguistic properties of sign languages, resulting in higher average DTW errors compared to S-MotionGPT. In contrast, our method sets a new state-of-the-art across all three benchmarks, achieving significantly lower DTW errors of 2.35, 1.71, and 1.38 on How2Sign, CSL-Daily, and Phoenix-2014T, respectively, using a single unified model.
4.2 Ablation Study
Figure 6: Qualitative ablation study for retrieval-enhanced SLG. (Left: How2Sign; Middle: CSL-Daily; Right: Phoenix-2014T.)
H2S CSL Ph-T Generation Reconstruction
@@ 4.2 Ablation Study
Body Hand Body Hand Body Hand Body Hand Body Hand Body Hand
7.92 3.07 21.52 7.74
7.11 2.63 6.79 2.14 19.88 6.97 24.29 5.83
6.82 2.35 6.24 1.71 4.77 1.38 19.37 6.65 23.52 5.13 25.79 6.78
Table 3: Study on the scalability of our approach. We use procrustes-aligned mean per joint position error (PA-MPJPE) [39, 78] to assess reconstruction performance of the decoupled tokenizer. (Ph-T: Phoenix-2014T.)
@@
Body Hand Body Hand Body Hand
Whole-body × × 9.49 3.56 9.22 3.35 8.03 3.23
Body Hands × 7.99 3.13 6.93 2.85 6.24 2.18
Body L-hand R-hand 6.82 2.35 6.24 1.71 4.77 1.38
Table 4: Ablation study on the decoupled tokenizer.
Decoding Method. We compare the performance of different decoding methods in Table 2. The traditional sequential decoding method is widely used in existing motion generation approaches [42, 74]. However, since tokens from different body parts are flattened into a single sequence, the number of decoding steps is tripled, leading to high inference latency (3.26s per video on How2Sign). By using three separate LM decoders, parallel decoding predicts token sequences for each body part simultaneously. This formulation significantly reduces latency (1.39s per video on How2Sign) but results in higher DTW errors due to the lack of information fusion across part-wise motion tokens. Notably, our multi-head decoding method achieves a better trade-off between efficiency and information integration, maintaining comparable latency to parallel decoding while delivering even superior generation quality than sequential decoding. We attribute this to the weighting mechanism (soft gating): since the decoder inputs are a weighted average of all part-wise tokens, errors in a single token may have a reduced impact on the overall performance.
Sign Retrieval. By leveraging external sign dictionaries, we develop a retrieval-enhanced SLG approach that incorporates retrieved word-level signs as additional generation conditions. As shown in Table 2, this retrieval mechanism significantly enhances the precision of generated signs, reducing the average DTW errors by 19.9%/19.6%/23.9% on How2Sign, CSL-Daily, and Phoenix-2014T, respectively. Furthermore, we perform a qualitative comparison in Figure 6. The figure demonstrates that the retrieved signs effectively guide the model’s generation: similar pose patterns from the retrieved signs, particularly in hand shapes, appear in the generated outputs. This guidance substantially improves the precision of the generated signs.
Scalability with Multilingual Data. Results in Table 3 show that incorporating multilingual data significantly improves generation performance, demonstrating the scalability of our approach. This improvement stems from two key factors: 1) the tokenizer benefits from more pose training data (right half of Table 3), which subsequently enhances the generator’s performance; 2) the presence of cross-lingual signs, as revealed in [69]. For example, the sign for “cold” in CSL and DGS shares similar body movements. Training on multilingual data helps the model better understand and generalize across these cross-lingual signs.
Decoupled Tokenizer. Inspired by the multi-cue characteristics of sign languages, we introduce a decoupled tokenizer that discretizes sign motion sequences for different body parts. In Table 4, we present an ablation study on various configurations: a single VQ-VAE for whole-body motions and one VAE for the body and another for both hands. The findings indicate that progressively separating the body and hands leads to consistently improved performance.